OpenStudy (anonymous):

sin 10 * sin 50 * sin 60 * sin70 =?

OpenStudy (anonymous):

@satellite73 @lgbasallote @mukushla @Callisto @Hero @hartnn @experimentX

hero (hero):

Hint: Calculator

OpenStudy (mimi_x3):

its not that easy..

hero (hero):

Yes it is. I plugged it into my calculator and got the answer

OpenStudy (mimi_x3):

It got to do with some identity i forgot the name of it :/

OpenStudy (anonymous):

@Hero no....calculator is allowed

hero (hero):

Unless those stars mean something other than multiply

OpenStudy (anonymous):

* === Multiplication.

hero (hero):

Well, sin(60) is easy

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=sin+10+*+sin+50+*+sin+60+*+sin70+ doesn't have a nice value.

OpenStudy (anonymous):

hm....hm...

OpenStudy (anonymous):

sin 10 sin (60 - 10) sin(60+10)

OpenStudy (anonymous):

sin 10 * (sin^2 60 - sin^10)

OpenStudy (anonymous):

sin 10 * (sin^2 60 - sin^2 10)

OpenStudy (anonymous):

i remember this, but i forgot how to do it some addition angle formula

OpenStudy (anonymous):

but nw wat ,,,,to do...

OpenStudy (experimentx):

what's the point ... you are getting it all in sin 10

OpenStudy (anonymous):

just plug the values for sin60

OpenStudy (experimentx):

probably you should plugin the value of sin 10

OpenStudy (anonymous):

is it \(\frac{1}{8}\)?

OpenStudy (anonymous):

nop...

OpenStudy (anonymous):

Yup...@satellite73 u r correct

OpenStudy (anonymous):

@satellite73 hw u got it..???

OpenStudy (anonymous):

memory

OpenStudy (experimentx):

hold on .. i guess I ignored the lower part. http://www.wolframalpha.com/input/?i=sin+10+*+sin+50+*+sin+60+*+sin70+ \[ \sqrt 3 \over 16\]

OpenStudy (experimentx):

yep .. plugin the value of sin(10)

OpenStudy (anonymous):

some trick of addition angle formulas will give you \[\frac{1}{8}+\sin(10)-\sin(10)\] i think

OpenStudy (anonymous):

Can u show the steps..

OpenStudy (anonymous):

honestly i forget and it if is what @experimentX wrote, then i guess i was off by \(\sin(60)\)

OpenStudy (experimentx):

sin 10 * (sin^2 60 - sin^2 10) = sin (10) 3/4 - sin^3(10) = (sin 30)/4 = 1/8

OpenStudy (anonymous):

@experimentX i dint get

OpenStudy (anonymous):

Any other way....! @satellite73

OpenStudy (anonymous):

ok give me a minute or two (or five)

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

oh lord here we go we need \(\cos(x)-\cos(y)=2\sin(\frac{x+y}{2})\sin(\frac{x-y}{2})\) and lets ignore \(\sin(60)\) we know what that is, we can multiply at the end

OpenStudy (anonymous):

take sin 60 = √3 / 2 outside √3/2 * 2(2 sin10 * sin50 * sin 70)

OpenStudy (anonymous):

using 2sinA sinB formula

OpenStudy (anonymous):

√3/4 [(cos 40 - 1/2) sin 70]

OpenStudy (anonymous):

so \[\sin(10)\sin(50)\sin(70)=\frac{1}{2}(2\sin(10)\sin(50))\sin(70)\] \[=\frac{1}{2}(\cos(40)-\cos(60))\sin(70)\] \[=\frac{1}{2}(\cos(40)-\frac{1}{2})\sin(70)\] \[=\frac{1}{2}(\cos(40)\sin(70))-\frac{1}{4}\sin(70)\] \[=\frac{1}{4}(2\cos(40)\sin(70))-\frac{1}{4}\sin(70)\] \[=\frac{1}{4}(\sin(110)+\sin(30))-\frac{1}{4}\sin(70)\] \[=\frac{1}{4}(\sin(110)+\frac{1}{2})-\frac{1}{4}\sin(70)\] this is getting exhausting

OpenStudy (anonymous):

@satellite73 didn't you take sin 60 outside ?

OpenStudy (anonymous):

i am ignoring it, because know what it is

OpenStudy (anonymous):

we are almost done

OpenStudy (anonymous):

hm..

OpenStudy (anonymous):

@satellite73 wonderful approach......but it is ugly...)))

OpenStudy (anonymous):

\[=\frac{1}{4}(\sin(110)+\frac{1}{2})-\frac{1}{4}\sin(70)\] \[\frac{1}{4}\sin(110)+\frac{1}{8}-\frac{1}{4}\sin(70)\] and since \(\sin(110)=\sin(70)\) you get \(\frac{1}{8}\)

OpenStudy (anonymous):

now you can multiply by \(\frac{\sqrt{3}}{2}\) to get the answer to your question i am happy to see a shorter way (worked out, not with steps missing, that is not shorter)

OpenStudy (anonymous):

but....@satellite73 the 1 st method...was much more easy..)))

OpenStudy (anonymous):

thxxx

OpenStudy (experimentx):

sin 10 sin 60 sin (60 - 10) sin(60+10) sin 10 \sqrt 3/2 ( sin 60 cos 10 - cos 60 sin 10) ( sin 60 cos 10 + cos 60 sin 10) sin 10 \sqrt 3/2 ( 3/4 cos^2 10 -1/4 sin^2 10) sin 10 \sqrt 3/2 ( 3/4 (1 -sin^2 10) -1/4 sin^2 10) sin 10 \sqrt 3/2 ( 3/4 -1 sin^2 10) sin 10 \sqrt 3/2 ( 3/4 -1 sin^2 10) \sqrt 3/2 ( 3 sin 10 -4 sin^3 10)/4 \sqrt 3/2 (sin 3*10)/4 sqrt 3/ 16 <- same as WA got http://www.wolframalpha.com/input/?i=sin+10+*+sin+50+*+sin+60+*+sin70+